OpenStudy (anonymous):

Your cell phone provider charges a monthly fee of 19.50 for 200 minutes.You are also charged 0.25 per min over 200 mins.Last month,your bill was 29.50 A. Write an equation for the situation above.Let m represent the number of mins. B.solve the equation you wrote in part (a)

OpenStudy (anonymous):

@DDCamp

OpenStudy (anonymous):

@AravindG

hero (hero):

m = number of minutes over 200 19.50 + 0.25m = 29.50 Solve for m

OpenStudy (anonymous):

thank you so much :)

hero (hero):

I understand how these can be confusing. They're designed to be tricky.

hero (hero):

One trick that might help you in figuring out how to setup your equation is to not only name your variable but also name the equation. For example: The Cost of Last Month's Bill is represented by: 19.50 + 0.25m = 29.50 where m is the amount of minutes over 200.

OpenStudy (anonymous):

Is the answer supposed to be 5

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@AravindG is the answer 5

OpenStudy (aravindg):

Nope. Try again.

OpenStudy (anonymous):

Is it 40

OpenStudy (aravindg):

Are u guessing from options?

OpenStudy (anonymous):

Yes I don't know how to solve it I tried but I got 5 from the equation that her gave me

OpenStudy (anonymous):

That hero gave me

OpenStudy (aravindg):

19.50 + 0.25m = 29.50 Subtract 19.5 on both sides. What do u get?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

29.50-19.50 is 10

OpenStudy (aravindg):

Now divide both sides by 0.25

OpenStudy (aravindg):

For easier calculation u can take 0.25 as 1/4

OpenStudy (anonymous):

2.5

OpenStudy (aravindg):

0.25m=10 Dividing both sides by 0.25?

OpenStudy (aravindg):

Try again

OpenStudy (anonymous):

.025

OpenStudy (anonymous):

M

OpenStudy (aravindg):

No.

OpenStudy (aravindg):

\[\dfrac{0.25}{0.25}m=\dfrac{10}{0.25}\]

OpenStudy (anonymous):

40

OpenStudy (aravindg):

How 40?

OpenStudy (anonymous):

10\0.25 is 40 according to the calculator

OpenStudy (aravindg):

Good :)

OpenStudy (aravindg):

Yes 40 is the answer!

hero (hero):

Congratulations @Pandabear78963 for figuring out how to the correct result. For future reference, we prefer that you show the steps you took to get to the result you believe is correct rather than just asking if a particular value is correct to avoid the appearance of guessing. For example, instead of asking if 5 is correct, take the following approach: Say, I believe 5 is correct because...then list the steps you took to get 5. Or more preferably, you can just show your steps then conclude that a certain value is the correct answer.

OpenStudy (aravindg):

Exactly. No learning takes place while guessing. And it would not help you at all in a real exam. So better change towards good learning than looking for an answer.

OpenStudy (anonymous):

Thank you hero and aravinindg

OpenStudy (aravindg):

yw :)